OpenStudy (lasttccasey):

I can't seem to remember or find a reliable thread on this.. Does arcsin(sin(x))=x? and does sin(arcsin(x))=x?

OpenStudy (mertsj):

Not all the time because the range of arcsin is -pi/2to pi/2 so if x = 5pi/6, for example, arcsin(sin(5pi/6))=pi/6

OpenStudy (lasttccasey):

My equation is actually S = 58.3+32.5cos((Π(t))/6) and I have to get "t" by itself

OpenStudy (lasttccasey):

Basically I need to know how to get "t" out of cos(Πt/6). Would arccos(cos(Πt/6)) = Πt/6?

OpenStudy (mertsj):

\[\frac{S-58.3}{32.5}=\cos \frac{pi}{6}t\] \[\frac{pi t}{6}=\cos^{-1} (\frac{S-58.3}{32.5})\] \[t=\frac{6\cos^{-1}( .0307692(S-58.3))}{pi}\]

OpenStudy (lasttccasey):

Interesting, I tried that the first time on my own but did a simple test with the number 5,\[x=\cos(5)\] I got x = .283662 when I reversed it it wouldn't work.. \[y=\arccos(x)\] y ended up being 1.28319. I now realize my calculator was in radians which I didn't even think about it.. needless to say I basically entered a degree but when I use Π/3 for example, it works fine. Sorry to waste your time, but thanks for the help anyways. At least I can solve this problem now.

OpenStudy (mertsj):

yw